OpenStudy (anonymous):

A total of 496 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was three times the number of adult tickets sold . How many adult tickets were sold?

Parth (parthkohli):

\[3x + x = 496\] \[4x = 496\] \[x = 124\] So, 124 student tickets and 124 * 3 = 372 adult tickets were sold.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

x=number of adult tx x+3x=496 4x=496 x=124